# Discrete Differential Geometry for $C^{1,1}$ Hyperbolic Surfaces of Non-Constant Curvature

### v3

@@ 6 Branch Points and Singular Edges
Note that in fig. 9 and fig. 10, almost every node in each of the asymptotic complexes is “regular,” meaning that it is the point of incidence for 4 distinct quads. This is not necessarily true of the center point, which is the point of incidence for $2n$ quads when the Amsler surface is comprised of $2n$ sectors. Such irregular points are termed “branch points.” Physically, branch points allow for the type of buckling and subwrinkling which is observed, for example, in the growth of flower petals or leafy lettuce, and is not accounted for by the “smooth” theory. These are discussed and analyzed in detail by Shearman and Venkataramani [SV21]. In particular, they observe that (1) surfaces with branch points cannot result from $C^{2}$ isometric immersions; in fact, they cannot be approximated arbitrarily well (with respect to local bending energy) by surfaces with bounded curvature and no branch points, and (2) $C^{1,1}$ isometric immersions—which allow for branch points—can, somewhat counterintuitively, produce surfaces with significantly smaller elastic energy, despite appearing more ruffled and wrinkled to the naked eye. One reason for this is that elastic energy concentrates near singular edges, where the angle between the asymptotic lines approaches $\pi$. However, as a hyperbolic surface is evolved outward from boundary data using the Lelieuvre formulas, one can forestall the development of a singular edge by inserting branch points, and thus continue the evolution of the surface without concentrating energy. To do so, Shearman and Venkataramani [SV21] develop a method of “surgery,” whereby one strategically cuts out portions of a surface and replaces the missing section with an odd number of Amsler or Pseudo-Amsler type sectors.
@@
[…]
Figure 11: Beginning with an $I\times I$ asymptotic complex, we remove a square from the upper right, and replace it with $m$ copies of itself. Note that, as quad graphs, each of the parallelograms we have added are equivalent to the original square which was removed. Further note that having replaced the original square with an odd number of sectors, we can consistently label edges on the interior of the net so that each quadrilateral has two $u$-edges (red) which are opposite each other, and two $v$-edges (blue) which are opposite each other, and that the resulting quad-graph can be “checkered” just as the original could.
@@
One can then specify units of distance to initialize $\mathbf{r}^{(k)}$ along these lines, and initialize the normal fields along these lines exactly as described above for the initial Amsler surface. Having done so, one can solve the discrete non-constant curvature Lelieuvre equations on each sector exactly as described above. We demonstrate this procedure in fig. 12. Note that the original sector (fig. 12, left) has a singular edge where the surface curls over on itself. We excise a square containing this edge, specify new axes and new boundary data in the missing sector, (fig. 12, center), and fill in the missing sector with a collection of new (Pseudo-)Amsler sectors (fig. 12, right).
This procedure can then be recursively applied to insert new branch points by excising pieces of the new (Pseudo-)Amsler sectors. The astute reader will note that the new sectors introduced through the branch point in fig. 12 still have a singular edge (or at least a very nearly singular edge). Inserting another branch point in each of the new sectors will ameliorate this. This is demonstrated in fig. 13, where we introduce another generation of branch points, and the newly introduced sectors do not have singular edges.
Figure 13: Left: The discrete hyperbolic surface from fig. 12 with additional branch points introduced in the new sectors. Right: The same discrete surface with 5 new sectors introduced at each branch point instead of 3.
While the demonstrations in fig. 12 and fig. 13 use only a single sector, these can pasted together as before to create closed surfaces whose normal field is continuous across the boundary lines. In fig. 14, we generate the same discrete surfaces as in fig. 9(d)-(f), but with branch points added in each sector. At the branch points, we alternately introduce $m=3$ or $m=5$ new sectors.
@@
Figure 14: The Amsler surfaces from fig. 9(d)-(f) with one generation of branch points introduced. Here the prescribed curvature is $K(p)=-(1+\varepsilon D(p))$, where $D(p)$ is geodesic distance to the origin. Each surface is comprised of $6$ Amsler sectors and we introduce $m=3$ or $m=5$ new sectors at the branch point.
@@ 7 Conclusion
Surfaces of non-constant negative curvature can serve as mathematical models for thin elastic sheets that arise in countless physical scenarios, both natural and artificial. In this paper, we describe a method for generating surfaces of prescribed non-constant negative curvature with distributed branch points. The key idea is to use the non-constant curvature version of the classical Lelieuvre formulas in both a continuous and discrete setting. In the case of non-constant curvature, the Lelieuvre formulas are no longer fully explicit, so we devise an iterative algorithm to numerically integrate them. We are especially interested in scenarios where curvature is a function of “material” coordinates, in particular the geodesic distance to some origin point. Accordingly, we describe a fast marching method for computing geodesic distance on a triangulated manifold. Using our method, we generate branched hyperbolic surfaces, which are comprised of bounded sectors patched together so as to maintain a continuous normal map throughout the surface. Finally, we demonstrate that our method is amenable to a sort of “surgery” to introduce branch points, whereupon sectors of a surface are excised and replaced by several new sectors. This allows for surfaces of non-constant negative curvature which exhibit the type of multi-generational buckling and wrinkling often observed in physical applications.
@@
There is a larger context for this work. Motivated by applications, there is interest in developing numerical methods for solving coupled systems of PDEs on (possibly evolving) surfaces. The typical approach has been through finite elements defined on a Lagrangian triangulation, i.e. a mesh that is fixed in the material, since physical processes on the surface are often naturally modeled through PDEs in material coordinates [BDNR12, DE13]. While successful, this approach does not exploit the particular simplifications that arise by having a grid that is adapted to the geometry, i.e. one that discretizes the asymptotic curves on the surface.
@@
We have considered a prototypical problem of this type, coupling “geometry”, namely the prescribed Gauss curvature equation, with “physics”, in this case the eikonal equation that determines the geodesic distance to the origin. To solve such coupled systems numerically, especially in scenarios involving potential branch points, we expect that a discretization that respects the extrinsic geometry of the surface and is “branch-point aware” will be more efficient. The underlying asymptotic complex for a hyperbolic surface does provide such a discretization and the mapping of this graph into the surface $\Omega$ gives a cellular decomposition into hyperbolic quadrilaterals. In particular, the sum of the angles in each quadrilateral is less than $2\pi$. This allows us to define a geometrically “natural” triangulation of $\Omega$ by bisecting quadrilaterals into 2 triangles, neither of which has an obtuse angle, thus generating a good-quality mesh. This procedure of obtaining a geometrically “natural” triangulation of $\Omega$ is illustrated in fig. 15. We expect that finite elements based on this approach will play an important role in solving coupled surface equations on evolving hyperbolic surfaces, especially in the presence of branch points.
@@
Figure 15: A discretization of surface like that in fig. 12 which respects both the “geometry” (by discretizing asymptotic curves) and the “physics” (by including a triangulation with no obtuse angles).
@@ Acknowledgments
The authors were supported in part by the Research Training Group in Data Driven Discovery at the University of Arizona (NSF award DMS-1937229) and by NSF award DMR-1923922. SV was supported in part by NSF award DMS-2108124.
@@ Appendix A Proof of Convergence of Algorithm 1 for Amsler-Type Surfaces
    Compute $(\mathbf{r}^{(n+1)}_{ij},\mathbf{N}^{(n+1)}_{ij})$ using (23) with curvature $K^{(n)}_{ij}$ (after having initialized the boundaries of $\mathbf{r}^{(n+1)}_{ij}$ using the same boundary data as $\mathbf{r}^{(0)}_{ij}$)
@@
$\mathbf{r}_{i+1,j}-\mathbf{r}_{ij}=\boldsymbol{\nu}_{i+1,j}\times\boldsymbol{\nu}_{ij}\,\,\,\,\,\text{ and }\,\,\,\,\,\mathbf{r}_{i,j+1}-\mathbf{r}_{ij}=-\boldsymbol{\nu}_{i,j+1}\times\boldsymbol{\nu}_{ij}$
@@
Note that $\mathbf{r}_{ij}$ and $\mathbf{r}_{ij}^{*}$ are given the same values on the discrete boundaries $i=0$ and $j=0$. We first prove that the normals do not differ by too much along the boundaries. We prove this for the difference $\|\mathbf{N}_{i,0}-\mathbf{N}_{i,0}^{*}\|$; the same bound will hold mutatis mutandis for $\|\mathbf{N}_{0,j}-\mathbf{N}_{0,j}^{*}\|$. Note that $\mathbf{N}_{i,0},\mathbf{N}_{i,0}^{*}$ are given by (33) with respective rotation angles $\delta_{i},\delta_{i}^{*}$ given by (34) using either $K$ or $K^{*}$. In what follows, all norms are 2-norms unless otherwise specified, and $C$ is a constant which changes from line to line and depends on some of the ambient data, but not upon $\|K-K^{*}\|_{\infty}$. First note that because $\arcsin$ has derivative bounded by $1$, we have
@@
$\|\mathbf{N}_{i,0}-\mathbf{N}_{i,0}^{*}\|$ =($=\left\|\left(\prod^{i-1}_{\ell=0}R_{\ell}\right)\mathbf{N}_{0,0}-\left(\prod^{i-1}_{\ell=0}R_{\ell}^{*}\right)\mathbf{N}_{0,0}\right\|$
=$\leq\left\|\prod^{i-1}_{\ell=0}R_{\ell}-\prod^{i-1}_{\ell=0}R_{\ell}^{*}\right\|$
=($=\left\|\left(\prod^{i-1}_{\ell=1}R_{\ell}\right)(R_{0}-R^{*}_{0})-\left(\prod^{i-1}_{\ell=1}R_{\ell}^{*}-\prod^{i-1}_{\ell=1}R_{\ell}\right)R^{*}_{0}\right\|$

### v4

@@ 6 Branch Points and Singular Edges
Note that in figure 8 and figure 9, almost every node in each of the asymptotic complexes is “regular,” meaning that it is the point of incidence for 4 distinct quads. This is not necessarily true of the center point, which is the point of incidence for $2n$ quads when the Amsler surface is comprised of $2n$ sectors. Such irregular points are termed “branch points.” Physically, branch points allow for the type of buckling and subwrinkling which is observed, for example, in the growth of flower petals or leafy lettuce, and is not accounted for by the “smooth” theory. These are discussed and analyzed in detail by Shearman and Venkataramani [53]. In particular, they observe that (1) surfaces with branch points cannot result from $C^{2}$ isometric immersions; in fact, they cannot be approximated arbitrarily well (with respect to local bending energy) by surfaces with bounded curvature and no branch points, and (2) $C^{1,1}$ isometric immersions—which allow for branch points—can, somewhat counterintuitively, produce surfaces with significantly smaller elastic energy, despite appearing more ruffled and wrinkled to the naked eye. One reason for this is that elastic energy concentrates near singular edges, where the angle between the asymptotic lines approaches $\pi$. However, as a hyperbolic surface is evolved outward from boundary data using the Lelieuvre formulas, one can forestall the development of a singular edge by inserting branch points, and thus continue the evolution of the surface without concentrating energy. To do so, Shearman and Venkataramani [53] develop a method of “surgery,” whereby one strategically cuts out portions of a surface and replaces the missing section with an odd number of Amsler or Pseudo-Amsler type sectors.
@@
We suppose that one has already generated a sector of an Amsler surface in the form of a discrete immersion $\{\mathbf{r}_{ij}\}_{0\leq i\leq I,0\leq j\leq J}$ using the method proposed in the previous section. For simplicity of exposition, we will assume that $I=J$, so that the immersion is defined on the nodes $Q=\{(i,j)\in\mathbb{Z}^{2}\,:\,0\leq i,j\leq I\}$. We further assume that the region to be modified also a square of the form $B=\{(i,j)\in\mathbb{Z}^{2}\,:b\leq i,j\leq I\}$ for some $b=1,\ldots,I$. We “delete” the subsurface described by $\{\mathbf{r}_{ij}\}_{b+1\leq i,j\leq I}$; this will be replaced with collection of new Amsler sectors given by immersions $\mathbf{r}^{(k)}:B\to\mathbb{R}^{3}$ and corresponding normal fields $\mathbf{N}^{(k)}:B\to\mathbb{S}^{2}$, for $k=1,\ldots,m$. In doing this, we are replacing the section $B$ of the asymptotic complex $Q$ with $m$ “copies” of itself. This is demonstrated with a $5\times 5$ asymptotic complex in figure 10. In the figure, we use $b=2$ and $m=3$, so that a $3\times 3$ square is removed and replaced with $3$ copies of itself. It is proven in [53], that in order for the resulting quad graph to still be an asymptotic complex, $m$ must be odd. Otherwise, there will be no consistent way to label edges so that each quadrilateral consists of two $u$-edges which are opposite each other, and two $v$-edges which are opposite each other. In figure 10, the pink-tiled parallelogram is the domain for $\mathbf{r}^{(1)}$ which shares a $u$-boundary with the original complex, the yellow-tiled parallelogram is the domain for $\mathbf{r}^{(2)}$, and the gray-tiled parallelogram is the domain for $\mathbf{r}^{(3)}$ which shares a $v$-boundary with the original parallelogram. We emphasize here that while the new parallelograms are equivalent as graphs to the excised square, the orientation of alternate parallelograms is flipped, in that the “right-hand” boundary is alternately comprised of $u$-edges or $v$-edges.
Figure 10: Beginning with an $I\times I$ asymptotic complex, we remove a square from the upper right, and replace it with $m$ copies of itself. Note that, as quad graphs, each of the parallelograms we have added are equivalent to the original square which was removed. Further note that having replaced the original square with an odd number of sectors, we can consistently label edges on the interior of the net so that each quadrilateral has two $u$-edges (red) which are opposite each other, and two $v$-edges (blue) which are opposite each other, and that the resulting quad-graph can be “checkered” just as the original could. $s$ and $\mathbf{N}$ are prescribed on nodes marked with a Green square, the closed circles mark nodes where $\mathbf{N}$ is specified.
@@
One can then specify units of distance to initialize $\mathbf{r}^{(k)}$ along these lines, and initialize the normal fields along these lines exactly as described above for the initial Amsler surface. Having done so, one can solve the discrete non-constant curvature Lelieuvre equations on each sector exactly as described above. We demonstrate this procedure in figure 11. Note that the original sector (figure 11, left) has a singular edge where the surface curls over on itself. We excise a square containing this edge, specify new axes and new boundary data in the missing sector, (figure 11, center), and fill in the missing sector with a collection of new (Pseudo-)Amsler sectors (figure 11, right).
This procedure can then be recursively applied to insert new branch points by excising pieces of the new (Pseudo-)Amsler sectors. The astute reader will note that the new sectors introduced through the branch point in figure 11 still have a singular edge (or at least a very nearly singular edge). Inserting another branch point in each of the new sectors will ameliorate this. This is demonstrated in figure 12, where we introduce another generation of branch points, and the newly introduced sectors do not have singular edges.
Figure 12: Left: The discrete hyperbolic surface from figure 11 with additional branch points introduced in the new sectors. Right: The same discrete surface with 5 new sectors introduced at each branch point instead of 3.
While the demonstrations in figure 11 and figure 12 use only a single sector, these can pasted together as before to create closed surfaces whose normal field is continuous across the boundary lines. In figure 13, we generate the same discrete surfaces as in figure 8(d)-(f), but with branch points added in each sector. At the branch points, we alternately introduce $m=3$ or $m=5$ new sectors.
@@
Figure 13: The Amsler surfaces from figure 8(d)-(f) with one generation of branch points introduced. Here the prescribed curvature is $K(p)=-(1+\varepsilon D(p))$, where $D(p)$ is geodesic distance to the origin. Each surface is comprised of $6$ Amsler sectors and we introduce $m=3$ or $m=5$ new sectors at the branch point.
7 The Strip Geometry: A Model for Edges
In the previous section, we considered a disk-like geometry where curvature depends on the geodesic distance from a central point, a model well-suited for objects that grow isotropically. However, many of the phenomena that motivate our work, such as the frilly edges of leaves or the crenellations on torn plastic sheets [51, 52], are better described by an asymptotically flat strip with a decorated edge. To model this, we construct an immersion using a composite curvature function that combines a power-law behavior near the edge with an asymptotically flat far-field.
7.1 Numerical Construction of the Strip Immersion
Our numerical construction of the strip immersion follows a three-step procedure. First, we define an asymptotic coordinate system using an analytically tractable seed surface. Second, we define our target intrinsic curvature model, which is a composite function designed to replicate experimental observations. Finally, we use the seed surface to provide boundary data for an iterative algorithm that deforms the initial surface until its geometry matches the target curvature.
We begin with a seed surface given by a Monge patch $(x,y,w(x,y))$ where $(x,y)\in\Omega\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(-\tfrac{\pi}{2},\tfrac{\pi}{2})\times[0,2]$ and $w(x,y)=Ae^{-y}\cos(x)$ with the constant $A$ to be specified later. This seed surface has Gauss curvature given by
$K=\frac{w_{xx}w_{yy}-w_{xy}^{2}}{(1+w_{x}^{2}+w_{y}^{2})^{2}}=\frac{-A^{2}e^{-2y}}{(1+A^{2}e^{-2y})^{2}}.$ (36)
In particular, the curvature depends only on $y$, is always negative $K(y)<0$, and approaches 0 as $y$ increases, so the surface is nearly flat near the “upper” boundary at $y=2$. Since $K<0$, we can foliate the surface by asymptotic curves that are given by
$\cos(x)\,dx^{2}+2\sin(x)\,dxdy+\cos(x)\,dy^{2}=0.$ (37)
Integrating gives two families of curves, from which we define the asymptotic coordinates $(u,v)$:
$\begin{aligned} u&=e^{-y}(1+\sin(x)),\\ v&=e^{-y}(1-\sin(x)),\end{aligned}\quad\quad\leftrightsquigarrow\quad\quad\begin{aligned} x&=\arcsin\left(\tfrac{u-v}{u+v}\right),\\ y&=-\log\left(\tfrac{u+v}{2}\right),\\ z&=A\sqrt{uv}.\end{aligned}$ (38)
The left and right boundaries $x=\pm\pi/2$ are themselves asymptotic curves where $u=0$ and $v=0$, respectively. A visualization of these coordinates on the seed surface is shown in Figure 14 (left and center panels).
This coordinate system suggests a computational domain equivalent to a truncated square grid, shown in Figure 14 (right panel). The discrete vertices $(i,j)$ with $i+j=M_{\text{min}}$ correspond to the line $y=2$, and those with $i+j=M_{\text{max}}$ correspond to $y=0$. This setup is ideal for a Goursat-type problem where initial data is specified on the “past” boundaries (black and green points) and the solution is evolved forward.
Figure 14: Left: 2D projection of asymptotic coordinates on the seed surface in $\Omega=[-\pi/2,\pi/2]\times[0,2]$. Center: The seed surface plotted in these coordinates. Right: The quadgraph of the computational domain. The bottom-left diagonal corresponds to the far boundary ($y=2$). The figure illustrates the Goursat problem setup: data for the normal $\mathbf{N}$ is specified on the boundary (black circles), while both $\mathbf{N}$ and the geodesic distance $s$ are specified at the corner (green squares).
While the seed surface provides a useful coordinate system, its curvature depends on the extrinsic coordinate $y$. To create a model based on the material’s internal geometry, we instead define an intrinsic curvature that depends on $s$, the geodesic distance from the wrinkled edge. We prescribe the total geodesic width of the strip to be 2, so the known far boundary corresponds to $s=2$.
To replicate the geometry of torn plastic sheets, we construct a composite function, $K_{\text{comp}}(s)$, that combines two distinct behaviors. Near the wrinkled edge, we use a power-law form, $K_{2}(s)\propto(s+\ell)^{-2}$, consistent with experimental observations [51, 52]. For the far-field, we adopt the asymptotically flat form from the seed surface’s curvature, $K_{1}(s)$, by replacing the role of extrinsic coordinate $y$ with the intrinsic distance $s$. By choosing parameters $A=e/\sqrt{2}$, $\ell=2$, and a contact point of $s_{0}=1$, these two forms are joined into the following $C^{1}$ function:
$K_{\text{comp}}(s)=\begin{cases}\frac{2}{(s+2)^{2}},&s<1\\ \frac{2\exp(2-2s)}{(2+\exp(2-2s))^{2}},&s\geq 1\end{cases}$ (39)
To find the immersion for this target curvature, we use the seed surface to provide boundary data and an initial guess, $\mathbf{r}^{(0)}_{ij},\mathbf{N}^{(0)}_{ij}$. In particular, the seed surface, with curvature $K_{1}(y)$ is very close to an immersion of surface with curvature $K_{1}(s)$ as the approximation $y\approx s$ is consistent in the far field. We then run Algorithm 1, with the modification that the fast marching method decreases $s$ starting with $s=2$ on the edge $y=2$, to update the immersion based on $K_{\text{comp}}(s)$, until the process converges. The result of this procedure, shown in Figure 15, is a surface that develops a singular cuspidal edge where the asymptotic curves from the $u$- and $v$- families touch tangentially.
Figure 15: Two views of the discrete surface with the intrinsic curvature $K_{\text{comp}}(s)$ given by (39). There is a singular edge where the $u$- and $v$-asymptotic curves touch tangentially and the surface is no longer the graph of a differentiable function.
As with the Amsler-type surfaces, the bending energy concentrates along this singular edge, which is not physically preferred. We therefore perform surgery to introduce branch points and forestall this singularity. The procedure is analogous to that of the previous section: we choose a cut point $b$, remove grid nodes $(i,j)$ with $i,j\geq b$, and replace the excised portion with $m$ copies, pasted along new boundary lines (Figure 16). The final branched surfaces, shown in Figure 17 for 3-fold, 5-fold and 7-fold subwrinkling have no singular edge. The surgically inserted regions successfully demonstrate the sort of high-frequency wrinkling observed at the edge of a torn plastic strip [51, 52].
Figure 16: Introducing a branch point via surgery removes the singular edge. The grid on the left represents the computational domain for the unbranched surface depicted in Fig. 15 that has a singular edge. On the right is a visualization of the surgery on the discrete computational domain, shown here for the case of 3-fold subwrinkling surface depicted in Fig. 17.
Figure 17: Surfaces with curvature given by $K_{\text{comp}}(s)$, where $s$ is the geodesic distance to the wrinkled edge. The far boundary is normalized to $s=2$. We have introduced branch points correponding to 3-fold (top), 5-fold (middle), and 7-fold (bottom) subwrinkling.
@@ 8 Conclusion
Surfaces of non-constant negative curvature can serve as mathematical models for thin elastic sheets that arise in countless physical scenarios, both natural and artificial. In this paper, we describe a method for generating surfaces of prescribed non-constant negative curvature with distributed branch points. The key idea is to use the non-constant curvature version of the classical Lelieuvre formulas in both a continuous and discrete setting. In the case of non-constant curvature, the Lelieuvre formulas are no longer fully explicit, so we devise an iterative algorithm to numerically integrate them. We are especially interested in scenarios where curvature is a function of “material” coordinates, in particular the geodesic distance to some point or edge. Accordingly, we describe a fast marching method for computing geodesic distance on a triangulated manifold. Using our method, we generate branched hyperbolic surfaces, which are comprised of bounded sectors patched together so as to maintain a continuous normal map throughout the surface. Finally, we demonstrate that our method is amenable to a sort of ‘surgery’ to introduce branch points, whereupon sectors of a surface are excised and replaced by several new sectors. This allows for surfaces of non-constant negative curvature which exhibit the type of multi-generational buckling and wrinkling often observed in physical applications.
@@
There is a larger context for this work. Motivated by applications, there is interest in developing numerical methods for solving coupled systems of PDEs on (possibly evolving) surfaces. The typical approach has been through finite elements defined on a Lagrangian triangulation, i.e. a mesh that is fixed in the material, since physical processes on the surface are often naturally modeled through PDEs in material coordinates [3, 16]. While successful, this approach does not exploit the particular simplifications that arise by having a grid that is adapted to the geometry, i.e. one that discretizes the asymptotic curves on the surface.
@@
We have considered a prototypical problem of this type, coupling “geometry”, namely the prescribed Gauss curvature equation, with “physics”, in this case the eikonal equation that determines the geodesic distance. To solve such coupled systems numerically, especially in scenarios involving potential branch points, we expect that a discretization that respects the extrinsic geometry of the surface and is “branch-point aware” will be more efficient. The underlying asymptotic complex for a hyperbolic surface does provide such a discretization and the mapping of this graph into the surface $\Omega$ gives a cellular decomposition into hyperbolic quadrilaterals. In particular, the sum of the angles in each quadrilateral is less than $2\pi$. This allows us to define a geometrically “natural” triangulation of $\Omega$ by bisecting quadrilaterals into 2 triangles, neither of which has an obtuse angle, thus generating a good-quality mesh. This procedure of obtaining a natural triangulation of $\Omega$ is illustrated in figure 18. We expect that finite elements based on this approach will play an important role in solving coupled surface equations on evolving hyperbolic surfaces, especially in the presence of branch points.
@@
Figure 18: A discretization of surface like that in figure 11 which respects both the “geometry” (by discretizing asymptotic curves) and the “physics” (by including a triangulation with no obtuse angles).
@@ Acknowledgments
The authors were supported in part by the Research Training Group in Data Driven Discovery at the University of Arizona (NSF award DMS-1937229) and by NSF award DMR-1923922. SV was supported in part by NSF awards DMS-2108124 and DMS- 2511503.
@@ Appendix A Proof of Convergence of Algorithm 1 for Amsler-Type Surfaces
[…]
@@
[…]
@@
Note that $\mathbf{r}_{ij}$ and $\mathbf{r}_{ij}^{*}$ are given the same values on the discrete boundaries $i=0$ and $j=0$. We first prove that the normals do not differ by too much along the boundaries. We prove this for the difference $\|\mathbf{N}_{i,0}-\mathbf{N}_{i,0}^{*}\|$; the same bound will hold mutatis mutandis for $\|\mathbf{N}_{0,j}-\mathbf{N}_{0,j}^{*}\|$. Note that $\mathbf{N}_{i,0},\mathbf{N}_{i,0}^{*}$ are given by (33) with respective rotation angles $\delta_{i},\delta_{i}^{*}$ given by (34) using either $K$ or $K^{*}$. In what follows, all norms are 2-norms unless otherwise specified, and $C$ is a constant which changes from line to line and depends on some of the ambient data, but not upon $\|K-K^{*}\|_{\infty}$. First note that because $\arcsin$ has derivative bounded by 11, we have
@@
$\|\mathbf{N}_{i,0}-\mathbf{N}_{i,0}^{*}\|$ $=\left\|\left(\prod^{i-1}_{\ell=0}R_{\ell}\right)\mathbf{N}_{0,0}-\left(\prod^{i-1}_{\ell=0}R_{\ell}^{*}\right)\mathbf{N}_{0,0}\right\|$
$\leq\left\|\prod^{i-1}_{\ell=0}R_{\ell}-\prod^{i-1}_{\ell=0}R_{\ell}^{*}\right\|$
$=\left\|\left(\prod^{i-1}_{\ell=1}R_{\ell}\right)(R_{0}-R^{*}_{0})-\left(\prod^{i-1}_{\ell=1}R_{\ell}^{*}-\prod^{i-1}_{\ell=1}R_{\ell}\right)R^{*}_{0}\right\|$